# On Efficient Variants of Segment Anything Model: A Survey

### v3

@@ 1 Introduction
The emergence of foundation models [1] has led to a thorough revolution in the field of artificial intelligence. Foundation models are large-scale neural networks pre-trained on massive data, which have powerful representing ability and strong generalization to perform various tasks. In the field of natural language processing, the recently popular research trend is about the large language models (LLMs) [2, 3], which has witnessed great development with prominent works like OpenAI’s GPT series [4, 5], Google’s PaLM series [6, 7] and Meta’s LLaMA series [8, 9]. Meanwhile, the success of Vision Transformer (ViT) [10] which firstly introduces the Transformer [11] architecture into the field of Computer Vision, has brought up a new era for vision foundation models (VFMs). Vision-language foundation models like CLIP [12], LLaVA [13], and Video-ChatGPT [14] etc., which aim at aligning the vision and language modalities, have demonstrated promising performance on plenty of downstream vision tasks[15, 16, 17, 18].
@@
Recently, a novel foundation model for general image segmentation, the Segment Anything Model (SAM), was proposed by Meta [19]. Being fully trained on their proposed SA-1B dataset, which consists of more than one billion masks and eleven million images, with the task of achieving valid segmentation given any prompt (e.g. points, boxes, masks and text), SAM is able to well generalize to a wide range of downstream tasks (e.g. edge detection, object proposal and instance segmentation) with proper prompts. Shortly after its emergence, SAM has garnered significant attention from the research community and results in a surge of related works exploring SAM’s generalization ability in various scenarios [20, 21, 22], including different image segmentation tasks [23, 24, 25, 26], video analysis tasks [27, 28, 29, 30] and 3D vision tasks [31, 32, 33]. With the huge success of SAM, the upgraded Segment Anything Model 2 (SAM 2) [34] is further proposed, aiming for efficient segmentation in both images and videos. SAM 2 introduces the streaming memory mechanism to extend SAM’s ability to videos. It is trained on both the SA-1B dataset and SA-V dataset which is their newly collected video segmentation dataset. As a result, SAM 2 naturally holds the powerful generalization and surpasses SAM when handling segmentation tasks [35, 36, 37, 38, 39, 40].
@@
Despite its success in a wide range of applications, the original Segment Anything Model, particularly SAM-H, faces significant limitations due to its slow runtime and high computational cost. These challenges become even more pronounced when SAM is deployed in resource-constrained or real-time environments. As the demand for deploying machine learning models in practical, resource-constrained scenarios continues to grow, SAM’s current design proves inefficient for widespread use. This has led to a pressing need for more lightweight and efficient variants that can maintain the model’s powerful segmentation capabilities while addressing these constraints. As the community strives to overcome these obstacles [41, 42, 43, 44, 45, 46, 47, 48], a comprehensive understanding of the latest advancements in making SAM more efficient is crucial. Consequently, it is both timely and necessary to conduct a detailed survey of the ongoing efforts aiming at improving SAM’s efficiency and extending its applicability across diverse environments.
@@
With the growing research related to SAM, several surveys have been presented to provide an overview from different perspectives [49, 50, 51, 52, 53, 54]. However, these existing surveys primarily focus on SAM’s downstream applications and exhibit several limitations: 1) None of them address the emerging field of improving SAM’s efficiency, which is gaining significant attention and is crucial for practical deployment in real-world applications. 2) With the exception of [53], most surveys lack a structured taxonomy that would allow for clearer organization and reference. 3) Most prior surveys focus on collecting and describing SAM-based methods but lack a systematic evaluation or comparison of these approaches. To address these gaps, we conduct this survey, not only to comprehensively review the development of efficient Segment Anything Models, but also to fairly evaluate and compare them. The main contributions of this work can be summarized as follows:
@@ 2 Preliminaries
In this section, we will first introduce the details and applications of the Segment Anything Model (SAM) in Section 2.1. In Section 2.2, we briefly review efficient backbones that could potentially replace SAM’s image encoder, and in Section 2.3, we discuss model compression techniques that could be applied to SAM. For a more comprehensive understanding of efficient backbones and model compression methods, we recommend referring to surveys [55, 56].
@@ 2.1 Segment Anything Model
Segment Anything Model [19] is a powerful foundation model in the image segmentation field, which can unify most image segmentation tasks by a fundamental segmentation task, i.e. the promptable segmentation task, with prompt engineering [57]. Another significant contribution of this project is the SA-1B dataset, which has over 1B masks from 11M licensed and privacy-preserving images. Trained with such abundant and high-quality data, SAM is expected to have strong robustness and generalization. The huge potential of SAM has quickly sparked researcher’s interest in both exploring its ability in a wide range of real-world applications and improving its architecture to segment more efficiently or accurately.
@@
Figure 1: The architectures of (a) SAM [19] and (b) the recent proposed SAM 2 [34].
Recently, Segment Anything Model 2 (SAM 2) [34] is proposed as a successor with a focus on efficient promptable visual segmentation (PVS) for both images and videos. To enable SAM 2 to segment anything in videos, researchers introduce the streaming memory mechanism into SAM’s original architecture. SAM 2 is trained from scratch by two stages: 1) Pre-training on SA-1B dataset with the promptable segmentation task for images; 2) Training on mixed data with the promptable segmentation task for images and videos. Similar to SAM, researchers build up a data engine to generate a large-scale dataset for video segmentation, named the SA-V dataset. With both manually annotated and automatically generated masklets (object segmentation in videos), SA-V finally collects 642.6K masklets across 50.9K videos. In this survey, we consider SAM 2 as an efficient SAM variant and include it in the evaluation and comparison.
@@ 2.1.1 Model
SAM consists of three components: an image encoder, a prompt decoder, and a mask decoder, as shown in Fig. 1 (a). The image encoder is an MAE [58] pre-trained Vision Transformer(ViT) with minimal modification. It takes the preprocessed images as input and outputs an image embedding for each image. The prompt decoder is to embed prompts: points, boxes, masks, and text. The two embeddings are then fed into the lightweight mask decoder, which is built upon two modified Transformer decoder blocks [11] and a few prediction heads, to generate valid masks. Based on SAM’s architecture, SAM 2 additionally introduces a streaming memory mechanism. Specifically, a memory encoder, a memory bank, and a memory attention module. The structure of SAM 2 is illustrated in Fig. 1 (b). With the memory mechanism, SAM 2 can handle videos frame by frame. The memory encoder generates the memory of the current frame and sends it to the memory bank. The memory bank stores memories of recent predictions, feature maps of prompted frames, and high-level semantic information of the target objects (i.e. object pointers). The memory attention mechanism aims to make the image embedding from the image encoder fully interact with the information from the memory bank, resulting in refined embedding. Apart from the memory mechanism, SAM 2 also adopts the MAE pre-trained Hiera [59] as the image encoder which is more efficient than ViT-H, with expectation for faster speed.
@@ 2.1.3 Application
As SAM and its successor SAM 2 have demonstrated strong generalization in plenty of zero-shot downstream tasks [19, 34], the community is diving into exploring their application in more scenarios.
@@
One of the major applications of SAM is medical image segmentation. According to [52], works in this area can be classified into two types. Some aim at testing SAM’s segmentation performance in CT images [23, 60], MRI images [61], pathological images [62], and so on. Others focus on improving the adaptation of SAM to these tasks by fine-tuning [20, 63], auto-prompting [64, 65, 66] or framework modification [67]. Furthermore, as these SAM-based works also suffered from high latency and heavy computational cost, improving medical SAM methods’ efficiency has increasingly gained attention in the community [68]. Several lightweight architectures [24, 69, 70, 71, 72] have emerged, achieving both high-speed and high-quality segmentation across various medical imaging modalities, most of which distilled MedSAM’s [20] image encoder into a lighter one to improve efficiency while maintaining the accuracy by subsequent fine-tuning. Optimization techniques during inference, for instance, exporting models to OpenVINO format and implementing an embedding caching mechanism [24, 69], could further speed up runtime efficiency, enabling variants to deploy on resource-constrained environments even without GPU.
@@
SAM is also applied to object detection across a diverse range of real-world scenarios [50, 73], including crack detection [25] and crater detection [74] in civil infrastructure defect assessment, crop disease and pest detection [75] in agriculture, anomaly detection [76] and remote sensing [77, 78, 79]. Moreover, Segment Anything has been adapted to Edit Everything [26], Inpaint Anything [80], and Caption Anything [81] to handle image editing tasks. Additionally, SAM is applied to provide priors for high-quality multi-modality image fusion [82]. Some researchers focus on exploring the uncertainty in SAM [83, 84]. For instance, UncertainSAM [83] has proposed effective uncertainty quantization for SAM, and UNICL-SAM[84] further leverage the uncertainty to guide segmentation.
@@
Apart from image segmentation tasks, SAM has been widely applied to various video tasks [53]. A large portion of the work concentrates on two basic tasks: video object segmentation (VOS) [27, 85, 86, 87] and video object tracking (VOT) [28, 88, 89]. Researchers also explore SAM’s application in tasks related to generation, for example, video super-resolution [29] and video dataset annotation generation [30, 90]. Besides these, SAM is further utilized as an intermediate tool in video editing tasks [91, 92]. Beyond 2D tasks, SAM is also extended to the 3D vision field. Shen et al. [31] applies SAM to 3D reconstruction and Dong et al. [32] applies it to 3D point cloud segmentation. The recent work [33] aims at achieving real-time segmentation for any 3D thing in an online setting. For the recently proposed SAM 2, some studies have already explored its application in both image and video tasks. A popular trend is to apply SAM 2 to medical images and video tasks [54]. Works like [35, 93, 36] evaluate SAM 2’s performance in medical images in both 2D and 3D modalities, while some others [38, 94] test it on surgical video segmentation tasks. Researchers are also seeking strategies to adapt SAM 2 to medical tasks better [95, 96]. Besides these, SAM 2 has been applied to some specific image segmentation tasks like digital pathology semantic segmentation [97], mesh part segmentation [98], solar panels segmentation [99], and 3D instance segmentation [100]. Moreover, several works [101, 102, 103, 39, 40] have improved SAM 2’s performance on the VOT task, the Referring Video Object Segmentation (RVOS) task and the challenging Large-scale Video Object Segmentation (LSVOS) task.
@@ 2.1.4 Limitation
Although SAM has demonstrated promising performance across various tasks, it still faces two key challenges in practical applications: 1) it often struggles to predict complete masks for fine structures, leading to imprecise boundaries; and 2) it is not real-time and remains resource-intensive, particularly when using heavy image encoders like ViT-H. To address these issues, some works aim to improve mask quality by utilizing high-resolution images [104, 105], integrating images and prompts [106], or introducing optimized prompts [107, 108], while others [41, 43, 46, 109] focus on creating more efficient architectures to reduce SAM’s time and resource consumption. Some recent works, such as  [110], have also started to find a better balance between model’s accuracy and efficiency. Previous surveys [20, 85, 31] have explored recent advances in enhancing SAM for higher-quality results. In this survey, we focus specifically on efforts to improve SAM’s efficiency.
Table 2: Parameters of SAM’s encoders. Results in this table refer to [43].
@@ 2.2.1 Efficient Vision Transformer
To reduce computational cost from a structural perspective, MobileViT [118], a pioneering hybrid architecture, creatively integrates CNN blocks (MobileNetV2 blocks [119]) with Transformer blocks into a single model. Subsequent works such as [114, 120, 116] basically follow this idea to build up efficient ViTs with hybrid structures, which have been widely used to substitute SAM’s heavy image encoder. In [41, 44], TinyViT [114] serves as the efficient backbone, while in [121] and [122], EfficientFormerV2 [120] and EfficientViT [116] supplant SAM’s original image encoder, respectively. Another influential ViT design, MetaFormer [123], which abstracts the attention mechanism into a broader concept called the token mixer, can deliver performance on par with Transformers using various token mixers. The simplest variant, PoolFormer, which uses pooling operations as the token mixer without introducing additional learnable parameters, has been leveraged as the base architecture to develop the light-weight image encoder for Lite-SAM [124]. The work [125] expedites Vision Transformer by utilizing the creatively proposed token clustering and token reconstruction mechanism, which has already been applied to SAM’s image encoder, resulting in the variant Expedit-SAM [126] with faster inference speed in image segmentation tasks.
Researchers have also made significant progress in optimizing the attention mechanism. It has been observed that the softmax operation within the attention mechanism contributes significantly to the overall computational cost. In EfficientViT [116], a novel ReLU linear attention is proposed to achieve a global receptive field with higher efficiency. This efficient backbone is further adopted in [122] to accelerate SAM. Improvements to the attention mechanism have also been made at the hardware level. FlashAttention [127] significantly reduces computation costs through techniques such as tiling, kernel fusion, and recomputation. And it is utilized in SAM acceleration works [128, 129] to reduce memory need and enhance computation efficiency.

### v4

@@ 1 Introduction
The emergence of foundation models [bommasani2021opportunities] has led to a thorough revolution in the field of artificial intelligence. Foundation models are large-scale neural networks pre-trained on massive data, which have powerful representing ability and strong generalization to perform various tasks. In the field of natural language processing, the recently popular research trend is about the large language models (LLMs) [zhao2023survey, wan2023efficient], which has witnessed great development with prominent works like OpenAI’s GPT series [NEURIPS2020_1457c0d6, achiam2023gpt], Google’s PaLM series [chowdhery2023palm, anil2023palm] and Meta’s LLaMA series [touvron2023llama, touvron2023llama2]. Meanwhile, the success of Vision Transformer (ViT) [dosovitskiy2020image] which firstly introduces the Transformer [NIPS2017_3f5ee243] architecture into the field of Computer Vision, has brought up a new era for vision foundation models (VFMs). Vision-language foundation models like CLIP [radford2021learning], LLaVA [liu2024llava], and Video-ChatGPT [maaz2023video] etc., which aim at aligning the vision and language modalities, have demonstrated promising performance on plenty of downstream vision tasks[li2024multimodal, gan2022vision, zou2024segment, tang2023video].
@@
Recently, a novel foundation model for general image segmentation, the Segment Anything Model (SAM), was proposed by Meta [kirillov2023segment]. Being fully trained on their proposed SA-1B dataset, which consists of more than one billion masks and eleven million images, with the task of achieving valid segmentation given any prompt (e.g. points, boxes, masks and text), SAM is able to well generalize to a wide range of downstream tasks (e.g. edge detection, object proposal and instance segmentation) with proper prompts. Shortly after its emergence, SAM has garnered significant attention from the research community and results in a surge of related works exploring SAM’s generalization ability in various scenarios [ma2024segment, chen2023sam, chen2023samfailssegmentanything], including different image segmentation tasks [roy2023sam, le2024medficientsam, ahmadi2023application, xie2023edit], video analysis tasks [zhang2023personalize, yang2023track, lu2023can, he2023scalable] and 3D vision tasks [shen2023anything, dong2023leveraging, xu2024embodiedsam]. With the huge success of SAM, the upgraded Segment Anything Model 2 (SAM 2) [ravi2024sam] is further proposed, aiming for efficient segmentation in both images and videos. SAM 2 introduces the streaming memory mechanism to extend SAM’s ability to videos. It is trained on both the SA-1B dataset and SA-V dataset which is their newly collected video segmentation dataset. As a result, SAM 2 naturally holds the powerful generalization and surpasses SAM when handling segmentation tasks [shen2024interactive, yamagishi2024zero, Wang_2025_CVPR, yu2024sam, tran20242nd, liu2024lsvos]. A further step in unified segmentation has been taken by the lately proposed Segment Anything Model 3 (SAM 3) [carion2025sam3segmentconcepts], which is built upon a image-level detector and a video-level tracker, with the goal of detecting, segmenting, and tracking all instances corresponding to the provided concept prompts in both images and videos.
@@
Despite its success in a wide range of applications, the original Segment Anything Model, particularly SAM-H, faces significant limitations due to its slow runtime and high computational cost. These challenges become even more pronounced when SAM is deployed in resource-constrained or real-time environments. As the demand for deploying machine learning models in practical, resource-constrained scenarios continues to grow, SAM’s current design proves inefficient for widespread use. This has led to a pressing need for more lightweight and efficient variants that can maintain the model’s powerful segmentation capabilities while addressing these constraints. As the community strives to overcome these obstacles [zhao2023fast, varadarajan2023squeezesam, zhang2023faster, shu2023tinysam, wang2024repvit, zhou2023edgesam, lv2024ptq4sam, chen20230], a comprehensive understanding of the latest advancements in making SAM more efficient is crucial. Consequently, it is both timely and necessary to conduct a detailed survey of the ongoing efforts aiming at improving SAM’s efficiency and extending its applicability across diverse environments.
@@
With the growing research related to SAM, several surveys have been presented to provide an overview from different perspectives [zhang2023survey, zhang2023comprehensive, zhang2023segment, zhang2024segment, zhang2024segment-videos, zhang2024unleashing]. However, these existing surveys primarily focus on SAM’s downstream applications and exhibit several limitations: 1) None of them address the emerging field of improving SAM’s efficiency, which is gaining significant attention and is crucial for practical deployment in real-world applications. 2) With the exception of [zhang2024segment-videos], most surveys lack a structured taxonomy that would allow for clearer organization and reference. 3) Most prior surveys focus on collecting and describing SAM-based methods but lack a systematic evaluation or comparison of these approaches. To address these gaps, we conduct this survey, not only to comprehensively review the development of efficient Segment Anything Models, but also to fairly evaluate and compare them. The main contributions of this work can be summarized as follows:
@@ 2 Preliminaries
In this section, we will first introduce the details and applications of the Segment Anything Model (SAM) in Section 2.1. In Section 2.2, we briefly review efficient backbones that could potentially replace SAM’s image encoder, and in Section 2.3, we discuss model compression techniques that could be applied to SAM. For a more comprehensive understanding of efficient backbones and model compression methods, we recommend referring to surveys [papa2024survey, xu2023survey].
@@ 2.1 Segment Anything Model
Segment Anything Model [kirillov2023segment] is a powerful foundation model in the image segmentation field, which can unify most image segmentation tasks by a fundamental segmentation task, i.e. the promptable segmentation task, with prompt engineering [liu2023pre]. Another significant contribution of this project is the SA-1B dataset, which has over 1B masks from 11M licensed and privacy-preserving images. Trained with such abundant and high-quality data, SAM is expected to have strong robustness and generalization. The huge potential of SAM has quickly sparked researcher’s interest in both exploring its ability in a wide range of real-world applications and improving its architecture to segment more efficiently or accurately.
@@
Figure 1: The architectures of (a) SAM [kirillov2023segment] and (b) the recent proposed SAM 2 [ravi2024sam].
Recently, Segment Anything Model 2 (SAM 2) [ravi2024sam] is proposed as a successor with a focus on efficient promptable visual segmentation (PVS) for both images and videos. To enable SAM 2 to segment anything in videos, researchers introduce the streaming memory mechanism into SAM’s original architecture. SAM 2 is trained from scratch by two stages: 1) Pre-training on SA-1B dataset with the promptable segmentation task for images; 2) Training on mixed data with the promptable segmentation task for images and videos. Similar to SAM, researchers build up a data engine to generate a large-scale dataset for video segmentation, named the SA-V dataset. With both manually annotated and automatically generated masklets (object segmentation in videos), SA-V finally collects 642.6K masklets across 50.9K videos. In this survey, we consider SAM 2 as an efficient SAM variant and include it in the evaluation and comparison.
@@ 2.1.1 Model
SAM consists of three components: an image encoder, a prompt decoder, and a mask decoder, as shown in Fig. 1 (a). The image encoder is an MAE [he2022masked] pre-trained Vision Transformer(ViT) with minimal modification. It takes the preprocessed images as input and outputs an image embedding for each image. The prompt decoder is to embed prompts: points, boxes, masks, and text. The two embeddings are then fed into the lightweight mask decoder, which is built upon two modified Transformer decoder blocks [NIPS2017_3f5ee243] and a few prediction heads, to generate valid masks. Based on SAM’s architecture, SAM 2 additionally introduces a streaming memory mechanism. Specifically, a memory encoder, a memory bank, and a memory attention module. The structure of SAM 2 is illustrated in Fig. 1 (b). With the memory mechanism, SAM 2 can handle videos frame by frame. The memory encoder generates the memory of the current frame and sends it to the memory bank. The memory bank stores memories of recent predictions, feature maps of prompted frames, and high-level semantic information of the target objects (i.e. object pointers). The memory attention mechanism aims to make the image embedding from the image encoder fully interact with the information from the memory bank, resulting in refined embedding. Apart from the memory mechanism, SAM 2 also adopts the MAE pre-trained Hiera [ryali2023hiera] as the image encoder which is more efficient than ViT-H, with expectation for faster speed.
@@ 2.1.3 Application
As SAM and its successor SAM 2 have demonstrated strong generalization in plenty of zero-shot downstream tasks [kirillov2023segment, ravi2024sam], the community is diving into exploring their application in more scenarios.
@@
One of the major applications of SAM is medical image segmentation. According to [zhang2024segment], works in this area can be classified into two types. Some aim at testing SAM’s segmentation performance in CT images [roy2023sam, hu2023sam], MRI images [mohapatra2023sam], pathological images [deng2023segment], and so on. Others focus on improving the adaptation of SAM to these tasks by fine-tuning [ma2024segment, li2024polyp], auto-prompting [hu2023skinsam, shaharabany2023autosam, Konwer_2025_CVPR] or framework modification [zhang2023semisam]. Furthermore, as these SAM-based works also suffered from high latency and heavy computational cost, improving medical SAM methods’ efficiency has increasingly gained attention in the community [ma2024efficient]. Several lightweight architectures [le2024medficientsam, pfefferle2024daft, wei2024rep, gao2024swin, hu2025breastlightsam] have emerged, achieving both high-speed and high-quality segmentation across various medical imaging modalities, most of which distilled MedSAM’s [ma2024segment] image encoder into a lighter one to improve efficiency while maintaining the accuracy by subsequent fine-tuning. Optimization techniques during inference, for instance, exporting models to OpenVINO format and implementing an embedding caching mechanism [le2024medficientsam, pfefferle2024daft], could further speed up runtime efficiency, enabling variants to deploy on resource-constrained environments even without GPU.
@@
SAM is also applied to object detection across a diverse range of real-world scenarios [zhang2023comprehensive, ji2024segment], including crack detection [ahmadi2023application] and crater detection [giannakis2023deep] in civil infrastructure defect assessment, crop disease and pest detection [li2023enhancing] in agriculture, anomaly detection [cao2023segment] and remote sensing [osco2023segment, wang2024samrs, Shan_2025_CVPR]. Moreover, Segment Anything has been adapted to Edit Everything [xie2023edit], Inpaint Anything [yu2023inpaint], and Caption Anything [wang2023caption] to handle image editing tasks. Additionally, SAM is applied to provide priors for high-quality multi-modality image fusion [Wu_2025_CVPR]. Some researchers focus on exploring the uncertainty in SAM [kaiser2025uncertainsam, Sheng_2025_CVPR]. For instance, UncertainSAM [kaiser2025uncertainsam] has proposed effective uncertainty quantization for SAM, and UNICL-SAM[Sheng_2025_CVPR] further leverage the uncertainty to guide segmentation.
@@
Apart from image segmentation tasks, SAM has been widely applied to various video tasks [zhang2024segment-videos]. A large portion of the work concentrates on two basic tasks: video object segmentation (VOS) [zhang2023personalize, zhang2023uvosam, cheng2023tracking, Mei_2025_CVPR] and video object tracking (VOT) [yang2023track, cheng2023segment, rajivc2023segment]. Researchers also explore SAM’s application in tasks related to generation, for example, video super-resolution [lu2023can] and video dataset annotation generation [he2023scalable, delatolas2024learning]. Besides these, SAM is further utilized as an intermediate tool in video editing tasks [zhao2023make, wu2023cvpr]. Beyond 2D tasks, SAM is also extended to the 3D vision field. Shen et al. [shen2023anything] applies SAM to 3D reconstruction and Dong et al. [dong2023leveraging] applies it to 3D point cloud segmentation. The recent work [xu2024embodiedsam] aims at achieving real-time segmentation for any 3D thing in an online setting.
For the subsequently proposed SAM 2, some studies have already explored its application in both image and video tasks. A popular trend is to apply SAM 2 to medical images and video tasks [zhang2024unleashing]. Works like [shen2024interactive, dong2024segment, yamagishi2024zero] evaluate SAM 2’s performance in medical images in both 2D and 3D modalities, while some others [yu2024sam, lou2025zero] test it on surgical video segmentation and works like [yin2025memory, xu2025tsms] attempt to improve SAM2’s accuracy on this task. Researchers are also seeking strategies to adapt SAM 2 to medical tasks better [yan2024biomedical, zhu2024medical]. Besides these, SAM 2 has been applied to some specific image segmentation tasks like digital pathology semantic segmentation [zhang2024sam2], mesh part segmentation [tang2024segment], solar panels segmentation [rafaeli2024prompt], and 3D instance segmentation [Zhao_2025_CVPR]. Moreover, several works [adamyan2025samrl, ding2024sam2long, Videnovic_2025_CVPR, Cuttano_2025_CVPR, tran20242nd, liu2024lsvos] have improved SAM 2’s performance on the VOT task, the Referring Video Object Segmentation (RVOS) task and the challenging Large-scale Video Object Segmentation (LSVOS) task. One popular trend is to optimize the memory mechanism, by capturing and integrating motion-aware [yang2024samurai, chen2025him2sam] or distractor-aware[Videnovic_2025_CVPR] information.
@@ 2.1.4 Limitation
Although SAM has demonstrated promising performance across various tasks, it still faces two key challenges in practical applications: 1) it often struggles to predict complete masks for fine structures, leading to imprecise boundaries; and 2) it is not real-time and remains resource-intensive, particularly when using heavy image encoders like ViT-H. To address these issues, some works aim to improve mask quality by utilizing high-resolution images [ke2024segment, huang2024hrsam], integrating images and prompts [Yu_2025_CVPR], or introducing optimized prompts [Radman_2025_CVPR, Liu_2025_CVPR], while others [zhao2023fast, zhang2023faster, zhou2023edgesam, xiong2024efficientsam] focus on creating more efficient architectures to reduce SAM’s time and resource consumption. Some recent works, such as  [wang2025tsda], have also started to find a better balance between model’s accuracy and efficiency. Previous surveys [ma2024segment, zhang2023uvosam, shen2023anything] have explored recent advances in enhancing SAM for higher-quality results. In this survey, we focus specifically on efforts to improve SAM’s efficiency.
Table 2: Parameters of SAM’s encoders. Results in this table refer to [zhang2023faster].
@@ 2.2.1 Efficient Vision Transformer
To reduce computational cost from a structural perspective, MobileViT [mehta2021mobilevit], a pioneering hybrid architecture, creatively integrates CNN blocks (MobileNetV2 blocks [sandler2018mobilenetv2]) with Transformer blocks into a single model. Subsequent works such as [wu2022tinyvit, li2023rethinking, cai2022efficientvit] basically follow this idea to build up efficient ViTs with hybrid structures, which have been widely used to substitute SAM’s heavy image encoder. In [zhao2023fast, shu2023tinysam], TinyViT [wu2022tinyvit] serves as the efficient backbone, while in [Zhao2023ESAM] and [zhang2024efficientvit], EfficientFormerV2 [li2023rethinking] and EfficientViT [cai2022efficientvit] supplant SAM’s original image encoder, respectively. Another influential ViT design, MetaFormer [yu2022metaformer], which abstracts the attention mechanism into a broader concept called the token mixer, can deliver performance on par with Transformers using various token mixers. The simplest variant, PoolFormer, which uses pooling operations as the token mixer without introducing additional learnable parameters, has been leveraged as the base architecture to develop the light-weight image encoder for Lite-SAM [fu2024lite]. The work [liang2022expediting] expedites Vision Transformer by utilizing the creatively proposed token clustering and token reconstruction mechanism, which has already been applied to SAM’s image encoder, resulting in the variant Expedit-SAM [ExpeditSAM] with faster inference speed in image segmentation tasks.
Researchers have also made significant progress in optimizing the attention mechanism. It has been observed that the softmax operation within the attention mechanism contributes significantly to the overall computational cost. In EfficientViT [cai2022efficientvit], a novel ReLU linear attention is proposed to achieve a global receptive field with higher efficiency. This efficient backbone is further adopted in [zhang2024efficientvit] to accelerate SAM. Improvements to the attention mechanism have also been made at the hardware level. FlashAttention [dao2022flashattention] significantly reduces computation costs through techniques such as tiling, kernel fusion, and recomputation. And it is utilized in SAM acceleration works [shen2024fastsam3d, songa2024sam] to reduce memory need and enhance computation efficiency.